. Suppose under TANF your state provides an income guarantee of $7,000 and a benefit reduction rate of 60%. The typical recipient can work up to 2,000 hours per year at a wage of $8 per hour. What would be the effective take-home wage rate if the recipient works 50 hours per year?
Answer #1

When you work for 50 hours per year, you earn $8*50 = $400. Your earnings are supplemented by $7,000 income guarantee with reduction in benefits at 60%. This implies that you will receive a total of 7000*40% = $2800. This makes your earnings worth $3200 for working 50 hours. Hence effective take home wage rate is 3200/50 = $64 per hour.